OpenStudy (anonymous):

Niki is building a kite. She has a rectangular piece of cloth that is 11 inches wide and 15 inches long. She needs to place a stick on the diagonal of the rectangle to make her kite. How long of a stick does she need? Round to the nearest tenth of an inch

OpenStudy (mathmale):

I'd suggest drawing a diagram and then considering what mathematical technique you'd use to find the length of the diagonal of the piece of cloth.

OpenStudy (zpupster):

 right triangle pythagorean theorem a^2+b^2=c^2 c is hypotenuse or stick 11^2 + 15^2 = c^2 121 + 225 = c^2 sqrt 336=c

OpenStudy (mathmale):

Great. Find the square root of 336 and round the result to the nearest 10th of an inch.

OpenStudy (zpupster):

yes

OpenStudy (mathmale):

Going to do that?

OpenStudy (anonymous):

thank you guys